OpenStudy (anonymous):

3. Given 4x – 8y = 8: a. Transform the equation into slope-intercept form. b. Find the slope and y-intercept of the line. c. Find the equation, in point-slope form, of the line that is perpendicular to this line and passes through the point (1, 2).

OpenStudy (anonymous):

@samsterz

OpenStudy (anonymous):

For the line with the equation below: a.Transform the equation into slope-intercept form. 4x–8y=8 8y=4x-8 y=x/2-1 .. b.Find the slope and y-intercept of the line. Standard form of equation of a line:y=mx+b, m=slope, b=y-intercept For given equation:4x–8y=8 slope=1/2 y-intercept=-1 .. c.Find the equation, in point-slope form, of the line that is perpendicular to this line and passes through the point (1, 2). slope of perpendicular line=-2 (negative reciprocal of slope of given line) point-slope form of equation: y1-y2=m(x1-x2) using coordinates of given point (1,2), Equation: y-2=-2(x-1)